OpenStudy (clamin):

PLEASE HELP!! MEDAL!! You put $5000 in a savings account which earns compound interest at a rate of 0.5% per month. If you dont make any deposits or withdrawls, how much will you have in your account in 30 years ( 360 months)? Round your answer to the nearest whole cent. a.$21,609.71 b.$5807.00 c.$22,324.85 d.$30,112.88

OpenStudy (anonymous):

Here is the compound interest formula: A=P(1+rn)nt A = Amount at maturity P = Principal Amount = $5,000 r = Annual interest rate in decimal = 12*.5 = 6%. In decimal: 0.06 n = compounding period (compounded how many times a year) = 12 t = years invested = 30 A=5,000(1+0.0612)12∗30=5,000(1.005)360=$30,112.88

OpenStudy (anonymous):

So your answer would be D.